OpenStudy (anonymous):

Gravity decreases with distance. which statement is also true? Orbital speed increases with distance. Orbital speed decreases with distance. Orbital speed remains constant with distance. Orbital speed varies based on the functions of friction.

OpenStudy (anonymous):

Orbital speed decreases with distance. That's one of the reasons the Moon takes longer to orbit the Earth than the International Space Station takes. The other reason is that the Moon has farther to go, since its orbit is larger.

OpenStudy (irishboy123):

\(m\omega^2 r = \dfrac{GMm}{r^2}\) \(\implies \left(\dfrac{2\pi}{T} \right)^2 = \dfrac{GM}{r^3}\) and \(\dfrac{mv^2}{ r} = \dfrac{GMm}{r^2}\) results \(\omega^2 \propto \dfrac{1}{r^3} \) \(T^2 \propto r^3 \) [Kepler's 3rd Law] \(v^2 \propto \dfrac{1}{r} \) slows down.....whatever way you look at it.